Your bank is offering you an APR of 17.6%. What is the monthly percentage rate you will pay if you don’t pay the credit card in full? Round to 2 decimal places.
1
Expert's answer
2020-11-04T16:45:30-0500

"EAR = [1 + (APR / m) ^ m] - 1=[(1+(\frac{0.176}{12})^{12}]-1=19.09"